OpenStudy (rainbow_rocks03):

What is the surface area of this rectangular prism? A. 120 cm2 B. 124 cm2 C. 240 cm2 D. 248 cm2

OpenStudy (anonymous):

The formula for the surface area of a rectangular prism is 2b+ph. just plug in the values and calculate.

OpenStudy (rainbow_rocks03):

I think the answer is C , am I correct? I got C because I did this...10*6*4=240

OpenStudy (anonymous):

That would be the volume, we are looking for the surface area. Did you see the formula i presented you?

OpenStudy (rainbow_rocks03):

Is the formula 2b+ph ? cause I don't know what that is...I have never seen that before.

OpenStudy (anonymous):

Yes, that is the formula. (2 * base area) + (perimeter * height)

OpenStudy (anonymous):

The formula is Surface Area = 2 (lwl+hl+hw)

OpenStudy (rainbow_rocks03):

So the base area would be 6?

OpenStudy (anonymous):

that is lw not lwl

OpenStudy (anonymous):

No, the base area would be 6*10.

OpenStudy (rainbow_rocks03):

6*10=60

OpenStudy (anonymous):

There are calculators you could use, for this questions. But S.A. = L.A. + 2B, so multiply perimeter times height, plus the base. :)

OpenStudy (anonymous):

And that's the base area. So 2(60) + perimeter * height.

OpenStudy (rainbow_rocks03):

Is the answer A cause I think it's A , Am I correct?

OpenStudy (anonymous):

@rainbow_rocks03 here is the formula: A=2(wl+hl+hw)

OpenStudy (anonymous):

No it would not be A. And @Famous5 that is not the only formula.

OpenStudy (rainbow_rocks03):

ALL OF YALL ARE CONFUSING ME.

OpenStudy (anonymous):

That is the easiest formula I know

OpenStudy (rainbow_rocks03):

I have never seen those formulas

OpenStudy (anonymous):

Well it is not the one we are currently using.

OpenStudy (anonymous):

What grade are you in rainbow_rocks03?

OpenStudy (anonymous):

because my formula is for 6th grade because I am only 6th grade

OpenStudy (rainbow_rocks03):

6th grade is the grade that I am in

OpenStudy (anonymous):

ok I am giving you a hint of what is the answer Hint: the answer is the greatest number

OpenStudy (rainbow_rocks03):

my answer is D

OpenStudy (anonymous):

correct

OpenStudy (rainbow_rocks03):

How is my answer 248 I want an explanation

OpenStudy (rainbow_rocks03):

No direct answer it is against code violation

OpenStudy (rainbow_rocks03):

@Famous5 HOW IS MY ANSWER 248

OpenStudy (anonymous):



OpenStudy (anonymous):

Im in 9th grade, and I gave you a formula that I remember from 8th grade. Its may seem hard but in reality it is much easier than what ever they taught you. The Surface Area of a solid is the sum of all of its surfaces. Maybe if you make a net of the prism it will be easier for you :)

OpenStudy (anonymous):

@LegendarySadist am I stealing your job?